OpenStudy (anonymous):

solve by the substitution method. 3x-5y=-10 x=-3-2y

jimthompson5910 (jim_thompson5910):

3x-5y=-10 3(-3-2y)-5y=-10 3(-3-2y)-5y=-10 -9-6y-5y=-10 -11y-9=-10 -11y=-10+9 -11y=-1 y=(-1)/(-11) y=1/11 ------------------------ Now that we know that y = 1/11, we can use it to find x x = -3-2y x = -3-2(1/11) x = -3-2/11 x = -33/11-2/11 x = -35/11 So the solutions are x = -35/11 and y = 1/11

OpenStudy (anonymous):

Thanks, are you a math teacher or just an algebra whiz?

jimthompson5910 (jim_thompson5910):

lol sorta both

jimthompson5910 (jim_thompson5910):

not yet a teacher

OpenStudy (anonymous):

good deal, your raking in the medals pal.

jimthompson5910 (jim_thompson5910):

lol thx

hero (hero):

He uses some program that allows him to enter solutions faster.

jimthompson5910 (jim_thompson5910):

yes and no

hero (hero):

Yes and yes

jimthompson5910 (jim_thompson5910):

sometimes i do, sometimes I don't

OpenStudy (anonymous):

hey what ever works, I say the person who helps me out the most gets the medals.

hero (hero):

It's the same program

hero (hero):

You use it everytime

OpenStudy (anonymous):

Hero, I appreciate your help also, really.

hero (hero):

and now he runs because he knows it's true

jimthompson5910 (jim_thompson5910):

half true

hero (hero):

I would buy it from you

OpenStudy (anonymous):

hey I want in on this program!haha

hero (hero):

Whatever it is. You claim that you use notepad ++ but I already had that and I don't see any advantage to using it

jimthompson5910 (jim_thompson5910):

well then you're not using it right

hero (hero):

Nope, and you haven't showed me how to use it right

jimthompson5910 (jim_thompson5910):

I didn't say I did or not

hero (hero):

You didn't

OpenStudy (anonymous):

hey it must take some kind of brains to dig up a program and use it to get these kind of answers. Float my boat dudes, I need all the help I can get!!!

hero (hero):

There he goes....running again. Just like last time

jimthompson5910 (jim_thompson5910):

lol this is pathetic....

OpenStudy (anonymous):

alrighty then, maths a bigger problem than this.....lol

jimthompson5910 (jim_thompson5910):

this is a non issue

OpenStudy (anonymous):

you bet

hero (hero):

Just show me how to do it

OpenStudy (anonymous):

here's a beer for both of you.

jimthompson5910 (jim_thompson5910):

how about you respect people more, then I'll think about it

hero (hero):

I don't want a beer. I want the program

jimthompson5910 (jim_thompson5910):

and actually help them out

hero (hero):

You weren't planning on showing me anyway. This proves it

OpenStudy (anonymous):

he might share if you don't squabble over petty stuff dude.

jimthompson5910 (jim_thompson5910):

yeah this is dumb and you know it

hero (hero):

Charlie, this convo is between me and jim now

OpenStudy (anonymous):

okay guys , it's a keg party.

jimthompson5910 (jim_thompson5910):

well it's charlies thread

hero (hero):

Jim, show me how to use it and I'll give you all the respect in the world

OpenStudy (anonymous):

sure is

jimthompson5910 (jim_thompson5910):

his rules

hero (hero):

scout's honor

jimthompson5910 (jim_thompson5910):

uh no

hero (hero):

See, I'd have to wait a lifetime for him to show me

hero (hero):

I think it's a pretty good deal.

hero (hero):

Jim is the type of person who creates companies and monopolizes them

jimthompson5910 (jim_thompson5910):

until you truly live up to your name

hero (hero):

See, you're only proving me right

jimthompson5910 (jim_thompson5910):

well if you say it's never going to happen, then it's just your fault

jimthompson5910 (jim_thompson5910):

not my problem

hero (hero):

You know I can't live up to my name. There's no such thing as Superman in real life

jimthompson5910 (jim_thompson5910):

metaphorically

hero (hero):

Whatever, I'll just figure it out myself

jimthompson5910 (jim_thompson5910):

you're obviously being selfish right now

hero (hero):

So are you

hero (hero):

Your selfishness is more selfish than mine

jimthompson5910 (jim_thompson5910):

whatever, this convo is over since no one is getting anywhere